Which of the subshell designations are possible and which are impossible? Possible Impossible Op 3p Id 3d 2f Answer Bank

pls help me with this assignment.
I think my arrangement might be mixed up
Answer #1

1. There are four quantam numbers- o Principal quantam number - denoted byn. It determines the main energy shell or level in② Magnetic quantam number (me)- It describes the behaviour of election in a magnetic field. for a given value of l, m. can haSince I can be no greater than noi n=1 = lman = 0 = 8 - and so on- (9) 3p = is possible - it has = n=3 b= (n-1) = (3-1) = 8 (